Find the volume of the solid generated by revolving about the x-axis bounded by the curve, y2 = 9x and the line y = 3x


1
Expert's answer
2021-06-08T07:08:02-0400

R.E.F image

"y=3x"

"y^\n2\n =ax ^\n2"

"ax=ax ^\n2"

"x=0,x=1"

Area bounded "=\u222b _\n0^\n1\n\u200b\n (3\\sqrt{ \nx}\n\u200b\n \u22123x)dx"

"=[ \\frac{32x ^{3\/2}}{3} \u2212 \\frac{3x^2}{2}] _0^\n1\n\u200b"

"=2− \frac{3}{2} =1/2 sq.units"